whats the linear system for
y=4x+6
y=-x+1

Found 2 solutions by Alan3354, sheldonbbtrocks:
Answer by Alan3354(69443)

whats the linear system for
y=4x+6
y=-x+1
============
That is a linear system.

Answer by sheldonbbtrocks(115)

All I can solve for is x, there is no y.
First,using the distributive property, multiply the second system by 4 to get the x's to cancel out:
4(-x+1)
4*-x=-4x
4*1=4
-4x+4
4x+6
-4x+4

Then we see that x=10